OpenStudy (anonymous):

Find the terminal point P(x, y) on the unit circle determined by t = (4/3)pi

OpenStudy (anonymous):

you are being asked to find \[(cos(\frac{4\pi}{3}),sin(\frac{4\pi}{3}))\] \[cos(\frac{4\pi}{3})=-\frac{1}{2}\] \[sin(\frac{4\pi}{3})=-\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

How do I know that those equal those fractions? Im sorry, I understand the unit circle slightly but I use the 30 60 90 and 45 45 90 triangles usually to solve for things. Do those triangles work in this situation as well? And how do I know if they are negative?